A 65-V generator supplies 3.3 kW of power.

How much current does the generator produce?
Answer #1

Solution:

Given:

Potential (V) = 65 V

Power (P) = 3.3 kW = 3300 W

Since: Power (P) = V I

Therefore: Current (I) = P / V = (3300 W) / (65 V) = 50.77 Amps

Hence: Current produced by the generator is 50.77 A.